You collect a sample of 188 randomly selected Wyomingites who married recently and ask them the cost of their wedding. The average amount is $34,575. According to the AWS the standard deviation is $15,009.

Provide a 87% confidence interval for the average cost of a wedding in Wyoming.
Use 2 decimal places.
Answer #1

Solutio,

Given that,

Point estimate = sample mean = = $ 34575

Population standard deviation =    = $ 15009

Sample size = n = 188

At 87% confidence level

= 1 - 87%  

= 1 - 0.87 =0.13

/2 = 0.065

Z/2 = Z0.065 = 1.514


Margin of error = E = Z/2 * ( /n)

= 1.514 * ( 15009 /  188 )

= 1657.29

At 87% confidence interval estimate of the population mean is,

  ± E

34575 ± 1657.29   

( $ 32917.71, $36232.29 )